OpenStudy (anonymous):

Find two consecutive even integers such that four times the first equals three times the second.

OpenStudy (xapproachesinfinity):

let 2n. 2n+2 be the consecutive even integers 4(2n)=3(2n+2)

OpenStudy (xapproachesinfinity):

that is the translation of the problem

OpenStudy (xapproachesinfinity):

find n and then multiply by 2 to find to get first integer what you get add 2 to it that gives the following integer

OpenStudy (anonymous):

2 mulip what again

OpenStudy (anonymous):

Two consecutive integers The first we can call n and to make sure it is even we double it, creating 2n the second has to be then next one because of the word consecutive so 2(n+1) Four times the first = 4(2n) Three times the second = 3 · 2(n+1) So 4(2n) = 3 · 2(n+1) now simplify 8n = 6n + 6 subtract 6n on both sides 2n = 6 n = 3 2(3) = 6 2(3+1) = 8 So the first number was 6 and the second was 8

OpenStudy (anonymous):

Ok how would i type that in as a answer

OpenStudy (anonymous):

The first consecutive integer: 6 The second consecutive integer: 8